### v2

# On the realization of infinite (continuous)

On the realization of infinite (continuous)
spin field representations in AdS4 space
I.L. Buchbinder1,2,    S.A. Fedoruk1,   A.P. Isaev1,3,   M.A. Podoinitsyn1
1Bogoliubov Laboratory of Theoretical Physics, Joint Institute for Nuclear Research,
141980 Dubna, Moscow Region, Russia
buchbinder@theor.jinr.ru, fedoruk@theor.jinr.ru, isaevap@theor.jinr.ru, mpod@theor.jinr.ru
2Center of Theoretical Physics, Tomsk State Pedagogical University,
634041 Tomsk, Russia
3Faculty of Physics, Lomonosov Moscow State University,
119991 Moscow, Russia
We study the symmetry properties of infinite spin fields in $\rm{AdS}_{4}$ space which are involved in the Lagrangian model proposed in arXiv:2403.14446 [hep-th] where the main role is played by operator constraints. It is shown that the conditions defining infinite spin states in $\rm{AdS}_{4}$ space are $\mathrm{SO}(2,3)$-invariant. It is found that in the model under consideration the Casimir operators are completely fixed by the constraint operators and only one of the Casimir operators is independent. It is shown that in this model, infinite spin fields in $\rm{AdS}_{4}$ space are described by the most degenerate representations of the $\mathrm{SO}(2,3)$ group.
PACS: 11.10.Ef, 11.30.-j, 11.30.Cp, 03.65.Pm, 02.40.Ky
Keywords: $\rm{AdS}_{4}$ space, Casimir operators, infinite spin representations
1 Introduction
One of the relatively recent trends in the higher spin field theory [^1] Current state of the higher spin field problem is considered, e.g., in the reviews [1, 2, 3]. is the study of the Lagrangian formulation of infinite (continuous) field models in $\rm{AdS}$ space. Various approaches to such a formulation are discussed in refs. [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. [^2] The review of the infinite (continuous) spin field models is given in ref. [14], for more recent references see e.g., [12]. In particular, a full classification of possible unitary and non-unitary continuous spin field models was given in the paper by R. Metsaev [9] within the light-cone formalism.
In a recent paper [12], a completely consistent covariant Lagrange formulation for bosonic infinite spin field in $\rm{AdS}_{4}$ space was constructed in the framework of the BRST approach to higher spin fields. The formulation is realized in the Fock state of the vectors $|{\Psi}\rangle$ which are annihilated by a set of operator constraints satisfying the closed algebra (see relations (2.4) bellow). The Lagrangian includes the real continuous parameter $\boldsymbol{\mu}$ and in the flat limit this Lagrangian is reduced to the known infinite spin field Lagrangian in Minkowski space in which $\boldsymbol{\mu}$ defines the corresponding Poincaré representation. This allows us to treat the vector $|{\Psi}\rangle$ as an infinite spin field in $\rm{AdS}_{4}$ space. Then natural questions arise as to what representation of the symmetry group of $\rm{AdS}$ space such a formulation corresponds and how it relates to Metsaev’s classification. In this paper we answer these questions. First, we show that the field $|{\Psi}\rangle$ in the paper [12] corresponds to the most degenerate continuous unitary representation of the Lie algebra $\mathfrak{so}(2,3)$. [^3] The notion of the most generate representation is based on the Gelfand theorem [15] which states that the number of independent invariant operators in the enveloping algebra acting in the Hilbert space of functions on the domain $X$ can be independent of the rank of the group. The most degenerate representation, if it exists in a given space, means that the number of these operators is one. The construction of the most degenerate continuous representations of the group $\mathrm{SO}(p,q)$ is discussed in the papers [16, 17, 18], see also [19]. Second, we show that the representations of the $\mathrm{SO}(2,3)$ group realized on the field $|{\Psi}\rangle$ correspond to the $\mathrm{SO}(2,3)$ representations in Metsaev’s classification. In this case, the eigenvalues of the $\mathfrak{so}(2,3)$ Casimir operators are determined by the dimensionless quantity $\boldsymbol{\mu}R$ where $R$ is the radius of $\rm{AdS}_{4}$ space.
The paper is organized as follows. In Section 2, we briefly review the basic results of the paper [12]. Here the infinite spin field in $\rm{AdS}_{4}$ space is defined, the operator constraints $L_{0},L,L^{+}$ are introduced, their algebra is written down and the flat limit is discussed. In Section 3, the geometric structure of $\rm{AdS}_{4}$ space is described in a form convenient for further consideration. In Section 4, we show that the operator constrains are the invariants of the $\mathrm{SO}(2,3)$ group and a more invariant operator $(\Pi)^{2}$ is introduced. In Section 5, we construct the Casimir operators $\mathcal{C}_{2}$ and $\mathcal{C}_{4}$ of the $\mathfrak{so}(2,3)$ algebra acting in the Fock space of fields $|{\Psi}\rangle$ under the conditions (2.4). In doing so, we have obtained two important statements. First, we proved that $\mathcal{C}_{2}$ and $\mathcal{C}_{4}$ are completely expressed in terms of the operators of the constraints (2.4) under consideration. Second, we show that only one of two these Casimir operators is independent. This exactly corresponds to the definition of the most degenerate representation for the $\mathrm{SO}(2,3)$ group. The summary is a collection of results.
2 Infinite spin field in $\rm{AdS_{4}}$ space
We begin with a brief summary of the basic results of the paper [12] on the Lagrangian formulation of the infinite spin field in $\rm{AdS_{4}}$ space.
The Lagrangian formulation under consideration was proposed within the BRST approach to higher spin field theory in terms of the Fock space with the vectors
$|{\Psi}\rangle=\sum\limits_{s_{1},s_{2}=0}^{\infty}\frac{1}{(s_{1}!s_{2}!)^{1/2}}\,\psi_{\alpha_{1}\ldots\alpha_{s_{1}}}^{\dot{\alpha}_{1}\ldots\dot{\alpha}_{s_{2}}}(x)c^{\alpha_{1}}\ldots c^{\alpha_{s_{1}}}\bar{c}_{\dot{\alpha}_{1}}\ldots\bar{c}_{\dot{\alpha}_{s_{2}}}|0\rangle\,,$ (2.1)
where $\psi_{\alpha_{1}\ldots\alpha_{s_{1}}}^{\dot{\alpha}_{1}\ldots\dot{\alpha}_{s_{2}}}(x)$ are the fields in $\rm{AdS_{4}}$ space with the coordinates $x^{\mu}$, $\mu=0,1,2,3$. Here $a_{\alpha}$, $\bar{a}^{\dot{\alpha}}$ and $c^{\alpha}=(\bar{a}^{\dot{\alpha}})^{\dagger}$, $\bar{c}_{\dot{\alpha}}=(a_{\alpha})^{\dagger}$ are the annihilation and creation operators, respectively. These operators satisfy the following commutation relations: [^4] The two-component Weyl spinor indices are raised and lowered by $\epsilon_{\alpha\beta}=-\epsilon^{\alpha\beta}$, $\epsilon_{\dot{\alpha}\dot{\beta}}=-\epsilon^{\dot{\alpha}\dot{\beta}}$: $\varphi^{\alpha}=\epsilon^{\alpha\beta}\varphi_{\beta}$, $\bar{\varphi}^{\dot{\alpha}}=\epsilon^{\dot{\alpha}\dot{\beta}}\bar{\varphi}_{\dot{\beta}}$. The relativistic Pauli matrices $\sigma^{m}_{\alpha\dot{\beta}}$ and $\tilde{\sigma}_{m}^{\dot{\alpha}\beta}$ satisfy the basic relation $\sigma_{m}\tilde{\sigma}_{n}+\sigma_{n}\tilde{\sigma}_{m}={}-2\eta_{mn}$ where $\eta_{mn}=\mathrm{diag}(-1,+1,+1,+1)$. We use the matrices $\sigma_{mn}=-\left(\sigma_{m}\tilde{\sigma}_{n}-\sigma_{n}\tilde{\sigma}_{m}\right)/4$, $\tilde{\sigma}_{mn}=-\left(\tilde{\sigma}_{m}\sigma_{n}-\tilde{\sigma}_{n}\sigma_{m}\right)/4$. Here and further the Latin indices take the values 0,1,2,3.
$[a_{\alpha},c^{\beta}]=\delta_{\alpha}^{\beta}\,,\qquad[\bar{a}^{\dot{\alpha}},\bar{c}_{\dot{\beta}}]=\delta_{\dot{\beta}}^{\dot{\alpha}}$ (2.2)
and act on the vacuum states $|0\rangle$ and $\langle 0|$, $\langle 0|0\rangle=1$ in the standard way:
$a_{\alpha}|0\rangle=\bar{a}^{\dot{\alpha}}|0\rangle=0\,,\qquad\langle 0|\,c^{\alpha}=\langle 0|\,\bar{c}_{\dot{\alpha}}=0\,.$ (2.3)
It was shown in [12] that the Lagrangian dynamics of the infinite spin field in $\rm{AdS_{4}}$ space is described by the vector (2.1) satisfying the conditions
$L_{0}|{\Psi}\rangle=0\,,\qquad L|{\Psi}\rangle=0\,,\qquad L^{+}|{\Psi}\rangle=0\,,\qquad U|{\Psi}\rangle=0\,,$ (2.4)
with the constraint operators
$L_{0}$ $=$ $(D)^{2}-\frac{1}{R^{2}}\left(N\bar{N}+N+\bar{N}\right)+\frac{2\boldsymbol{\mu}}{R}+\frac{1}{2R^{2}}\,K^{2}\,,$ (2.5)
$L$ $=$ $i\,(a\sigma^{m}\bar{a})D_{m}-\boldsymbol{\mu}-\frac{1}{4R}\,K^{2}\,,$ (2.6)
$L^{+}$ $=$ $i\,(c\sigma^{m}\bar{c})D_{m}-\boldsymbol{\mu}-\frac{1}{4R}\,K^{2}\,,$ (2.7)
$U$ $=$ $N-\bar{N}\,,$ (2.8)
where the quantities $N$, $\bar{N}$, $K$ are defined by the expressions
$N=c^{\alpha}a_{\alpha}\,,\qquad\bar{N}=\bar{c}_{\dot{\alpha}}\bar{a}^{\dot{\alpha}}\,,\qquad K=N+\bar{N}+2\,,$ (2.9)
the real parameter $\boldsymbol{\mu}$ is associated with the infinite spin representation, the positive constant $R$ is the curvature radius of $AdS_{4}$ space,  [^5] In [12], the parameter $\kappa$ was used, which is equal to $\kappa=-1/R^{2}$. and
$D_{m}=e_{m}^{\mu}D_{\mu}\,,\qquad D_{\mu}=\partial_{\mu}+\frac{1}{2}\,\omega_{\mu}{}^{kl}\mathcal{M}_{kl}$ (2.10)
is the covariant derivative. Here $e_{m}^{\mu}$ is the matrix inverse to the vierbein $e^{m}_{\mu}$, $e_{m}^{\mu}e^{n}_{\mu}=\delta_{n}^{m}$, $e^{m}_{\mu}e_{m\nu}=g_{\mu\nu}$, the real field $\omega_{\mu}{}^{kl}(x)$ is the spin-connection and the operators
$\mathcal{M}_{mn}=M_{mn}+\bar{M}_{mn}\,,\qquad M_{mn}:=c^{\alpha}\,(\sigma_{mn})^{\;\;\beta}_{\alpha}\,a_{\beta}\,,\quad\bar{M}_{mn}:=\bar{c}_{\dot{\alpha}}\,(\tilde{\sigma}_{mn})^{\dot{\alpha}}_{\;\;\dot{\beta}}\,\bar{a}^{\dot{\beta}}\,,$ (2.11)
are the anti-Hermitian generators, $(\mathcal{M}_{mn})^{\dagger}={}-\mathcal{M}_{mn}$, of the Lorentz algebra $\mathfrak{so}(1,3)$. In (2.5), the operator $(D)^{2}$ has the form
$(D)^{2}=D^{m}D_{m}+e^{\mu}_{m}\omega_{\mu}^{\;\;ml}D_{l}=\frac{1}{\sqrt{-g}}\,D_{\mu}\sqrt{-g}g^{\mu\nu}D_{\nu}\,,$ (2.12)
where $g=\det g_{\mu\nu}$. Relation (2.12) allows one to call the operator $(D)^{2}$ the generalized Laplace-Beltrami operator in the Fock space. Additionally, in (2.6) and (2.7), we use the non-indexed notation: $(c\sigma^{m}\bar{c}):=c^{\alpha}\sigma^{m}_{\alpha\dot{\alpha}}c^{\dot{\alpha}}$, $(a\sigma^{m}\bar{a}):=a^{\alpha}\sigma^{m}_{\alpha\dot{\alpha}}a^{\dot{\alpha}}$.
The operators (2.5)-(2.8) form a closed algebra in terms of commutators: the only nonzero commutator has the form
$[L^{+},L]=KL_{0}+\frac{1}{R}\,(K+1)L+\frac{1}{R}\,(K-1)L^{+}\,.$ (2.13)
Before going further, let us make a number of important comments about the defining conditions (2.4).
First, the operator $U$ in the last condition (2.4) is the $\mathrm{U}(1)$ symmetry generator. The solution of the equation $U|{\Psi}\rangle=0$ is the vector (2.1), where $s_{1}$ and $s_{2}$ are equal to one another, $s_{1}=s_{2}=s$, and the expansion contains only the fields $\psi_{\alpha_{1}\ldots\alpha_{s}}^{\dot{\alpha}_{1}\ldots\dot{\alpha}_{s}}(x)$ with the equal number of undotted and dotted spinor indices. These are exactly the states, resolving the constraints $U|{\Psi}\rangle=0$, that were used in constructing the Lagrangian formulation in [12].
Second, the limit $R\to\infty$ (together with $\omega_{\mu}{}^{kl}=0$ and $e_{m}^{\mu}=\delta_{m}^{\mu}$) is well defined: it simply switches off the external gravitational field. In this limit, equations (2.4) yield the conditions that define the well-known representation of the Poincáre group with infinite spin in flat space. Indeed, in the flat limit the operators (2.5)-(2.7) take the form
$\left.L_{0}\right|_{R\to\infty}=\partial^{m}\partial_{m}\,,\qquad\left.L\right|_{R\to\infty}=i\,(a\sigma^{m}\bar{a})\,\partial_{m}-\boldsymbol{\mu}\,,\qquad\left.L^{+}\right|_{R\to\infty}=i\,(c\sigma^{m}\bar{c})\,\partial_{m}-\boldsymbol{\mu}\,.$ (2.14)
In this case, the anti Hermitian generators of the translations are $P_{m}=\partial_{m}$ and the square of the Pauli-Lubanski pseudovector $W_{m}=-\frac{1}{2}\,\epsilon_{mnkl}P^{n}\mathcal{M}^{kl}$ is equal to
$W^{m}W_{m}\ ={}-(c\sigma^{m}\bar{c})(a\sigma^{n}\bar{a})\,\partial_{m}\partial_{n}\ +\ \left[\frac{N}{2}\left(\frac{N}{2}+1\right)+\frac{\bar{N}}{2}\left(\frac{\bar{N}}{2}+1\right)+\frac{N\bar{N}}{2}\right]\partial^{m}\partial_{m}\,.$ (2.15)
Taking into account relations (2.14), one gets that the condition $W^{m}W_{m}=\boldsymbol{\mu}^{2}$ holds for vectors obeying the first three equations (2.4) in the flat case $R\to\infty$. The last condition in (2.8) removes the degeneracy arising from the use of additional spinor variables in our approach.
Thus, in the flat case $R\to\infty$, the conditions (2.4) define the infinite spin field. On the other hand, equations (2.4) for a nonzero curvature $\kappa=-1/R^{2}\neq 0$ provide a consistent dynamical description of such a system in $AdS_{4}$ space.
Our main aim further is to clarify which representations of the $\rm{AdS_{4}}$ symmetry group correspond to the dynamical infinite spin field system constructed above.
3 Geometry of $\rm{AdS_{4}}$ space
One of the ways to define the anti-de Sitter space $\rm{AdS_{4}}$ of the radius $R$ is to identify it with the hypersurface
$X^{M}X^{N}\eta_{MN}=-R^{2}$ (3.1)
in five-dimensional two-time space $\mathbb{R}^{2,3}$ with the metric $\eta_{MN}=\mathrm{diag}(-,+,+,+,-)$ and the coordinates $X^{M}$, $M=0,1,2,3,4$. The manifold defined by the condition (3.1) is the coset $\mathrm{SO}(2,3)/\mathrm{SO}(1,3)$. The symmetry generators of $\rm{AdS_{4}}$ space are 666 Throughout this paper we use the anti-Hermitean form of generators: for example, $(J_{MN})^{\dagger}={}-J_{MN}$.
$J_{MN}=X_{M}\frac{\partial}{\partial X^{N}}-X_{N}\frac{\partial}{\partial X^{M}}\,,$ (3.2)
where $X_{M}=\eta_{MN}X^{N}$ form the $\mathfrak{so}(2,3)$ algebra:
$[J_{MN},J_{KL}]=\eta_{NK}J_{ML}+\eta_{ML}J_{NK}-\eta_{MK}J_{NL}-\eta_{NL}J_{MK}\,.$ (3.3)
For the generators $J_{mn}$, $m=0,1,2,3$ together with the generators
$P_{m}:=R^{-1}J_{m4}\,,$ (3.4)
the algebra (3.2) takes the form of the $\rm{AdS_{4}}$ symmetry algebra:
$\begin{array}[]{c}[P_{m},P_{n}]=R^{-2}J_{mn}\,,\\[7.0pt] [J_{mn},P_{l}]=\eta_{nl}P_{m}-\eta_{ml}P_{n}\,,\qquad[J_{mn},J_{kl}]=\eta_{nk}J_{ml}+\eta_{ml}J_{nk}-\eta_{nl}J_{mk}-\eta_{mk}J_{nl}\,,\end{array}$ (3.5)
where $\eta_{mn}=\mathrm{diag}(-,+,+,+)$ and the operators $J_{mn}$ form the $\mathfrak{so}(1,3)$ subalgebra.
Another way to describe this space is to use four independent coordinates $x^{\mu}$, $\mu=0,1,2,3$. Expressions for the coordinates $X^{M}$ in terms of the curved space coordinates $x^{\mu}$, resolving the condition (3.1), can be chosen in the form (see e.g., [20])
$X^{m}=e_{\mu}^{m}(x)\,x^{\mu}\,,\qquad X^{4}=R\left(\frac{2}{G(x)}-1\right)\,,$ (3.6)
where
$G(x)=1-\frac{\eta_{\mu\nu}x^{\mu}x^{\nu}}{4R^{2}}$ (3.7)
and
$e_{\mu}^{m}(x)=G^{-1}(x)\,\delta_{\mu}^{m}$ (3.8)
is the vierbein. The inverse matrix is formed by the vectors
$e^{\mu}_{m}(x)=G(x)\,\delta^{\mu}_{m}\,,$ (3.9)
whereas the metric tensors are
[…]
Let us emphasize that in (3.7) and (3.10) the tensors $\eta_{\mu\nu}=\eta^{\mu\nu}=\mathrm{diag}(-,+,+,+)$ are constant despite the fact that they have the curved space indices.
In the case where the $\rm{AdS_{4}}$ metric quantities have the form (3.7)-(3.10), the spin connection is written as follows:
$\omega_{m}{}^{kl}(x)=e^{\mu}_{m}\omega_{\mu}{}^{kl}(x)=R^{-2}\,{\delta_{m}^{[k}\delta_{\nu}^{l]}x^{\nu}}\,.$ (3.11)
At the same time, we obtain the standard form of the $\rm{AdS_{4}}$ curvature tensor:
$\mathcal{R}_{mn}{}^{kl}=e^{\mu}_{m}e^{\nu}_{n}\mathcal{R}_{\mu\nu}{}^{kl}(x)={}-R^{-2}\left(\delta_{m}^{k}\delta_{n}^{l}-\delta_{m}^{l}\delta_{n}^{k}\right)\,.$ (3.12)
In the curve variable $x^{\mu}$ in $\rm{AdS_{4}}$ space (3.1) the generators $J_{mn}$ defined in (3.2) are represented only by orbital generators $J_{mn}=\mathcal{L}_{mn}$ which have the form (here $\eta_{m\mu}=\mathrm{diag}(-,+,+,+)$):
$\mathcal{L}_{mn}=\left(\eta_{m\mu}\delta_{n}^{\nu}-\eta_{n\mu}\delta_{m}^{\nu}\right)x^{\mu}\,\frac{\partial}{\partial x^{\nu}}\,.$ (3.13)
In this case, the generators (3.4) take the form
$P_{m}=\left(1+\frac{\eta_{\nu\lambda}x^{\nu}x^{\lambda}}{4R^{2}}\right)\delta^{\mu}_{m}\partial_{\mu}-\frac{1}{2R^{2}}\,\eta_{m\mu}x^{\mu}x^{\lambda}\partial_{\lambda}\,.$ (3.14)
This expression (3.14) is represented as
$P_{m}=e^{\mu}_{m}\left(\partial_{\mu}-\frac{1}{2}\,\omega_{\mu}{}^{kl}\mathcal{L}_{kl}\right),$ (3.15)
where $e^{\mu}_{m}$, $\omega_{\mu}{}^{kl}$ and $\mathcal{L}_{kl}$ are defined in (3.9), (3.11) and (3.13), respectively.
The operators (3.15) and (3.13) realize the representation of the $\rm{AdS_{4}}$ algebra (3.5) in the space of scalar functions depending on the curve coordinate $x^{\mu}$. However, the infinite spin system under consideration is defined by conditions (2.4) on the Fock space vector (2.2), (2.3), where the creation $c^{\alpha}$, $\bar{c}_{\dot{\alpha}}$ and annihilation $a_{\alpha}$, $\bar{a}^{\dot{\alpha}}$ operators act, carrying indices of the tangent space structural group $\mathrm{SO}(1,3)$. This additional structure is taken into account by the replacement $\mathcal{L}_{mn}\ \to\ \mathcal{L}_{mn}+\mathcal{M}_{mn}$ in the operators (3.15) and (3.13) where $\mathcal{M}_{mn}$ are defined in (2.11).
Thus, as generators of the $\mathrm{SO}(1,3)$ symmetry of the system under study, we consider the operators
$J_{mn}=\mathcal{L}_{mn}+\mathcal{M}_{mn}\,,$ (3.16)
where $\mathcal{L}_{mn}$ and $\mathcal{M}_{mn}$ are defined in (3.13) and (2.11), respectively. Four residual generators are
$P_{m}=e^{\mu}_{m}\left(\partial_{\mu}-\frac{1}{2}\,\omega_{\mu}{}^{kl}J_{kl}\right),$ (3.17)
where $J_{kl}$ are defined in (3.16). It is easy to check by direct calculations that the operators (3.16) and (3.17) obey relations (3.5) and thus form the $\mathfrak{so}(2,3)$ algebra.
Note that some of the commutators in (3.5) can be represented as
$[P_{n},P_{m}]={}-\frac{1}{2}\,\mathcal{R}_{nm}{}^{kl}J_{kl}\,,$ (3.18)
where $\mathcal{R}_{nm}{}^{kl}$ are defined in (3.12).
4 $\mathrm{SO}(2,3)$ symmetry of infinite spin field
Having found expressions for the $\mathfrak{so}(2,3)$ symmetry generators, we now analyze the $\mathrm{SO}(2,3)$ symmetry of the conditions (2.4) that determine the infinite spin field in $\rm{AdS_{4}}$ space. The operator constraints in relations (2.5)-(2.8) are expressed in terms of the operator (2.10), which is written as follows:
$D_{m}=e_{m}+\frac{1}{2}\,\omega_{m}{}^{kl}\mathcal{M}_{kl}\,,$ (4.1)
where $\omega_{m}{}^{kl}=e^{\mu}_{m}\omega_{\mu}{}^{kl}$ and the first term equals
$e_{m}:=e_{m}^{\mu}\partial_{\mu}\,.$ (4.2)
The operators (4.2) form the algebra
$[e_{m},e_{n}]=\mathcal{E}_{mn}{}^{l}e_{l}\,,$ (4.3)
where
$\mathcal{E}_{mn}{}^{l}:=(e_{m}e_{n}^{\mu}-e_{n}e_{m}^{\mu})\,e_{\mu}^{l}$ (4.4)
is the anholonomy tensor, which is equal to $\mathcal{E}_{mn}{}^{l}=-x^{\mu}\eta_{\mu[m}\delta_{n]}^{l}/R^{2}$ in $\rm{AdS_{4}}$ space (3.8), (3.9). In these terms, the commutators of the operators (4.1) are written as follows:
$[D_{m},D_{n}]=\mathcal{E}_{mn}{}^{l}D_{l}+\frac{1}{2}\,\mathcal{R}_{mn}{}^{kl}\mathcal{M}_{kl}\,,$ (4.5)
whereas the commutation relations of the operators (4.1) with the generators (3.16) and (3.17) of the $\rm{AdS_{4}}$ algebra $\mathfrak{so}(2,3)$ are
$[P_{m},D_{k}]=\omega_{mk}{}^{l}D_{l}\,,\qquad[J_{mn},D_{k}]=\eta_{nk}D_{m}-\eta_{mk}D_{n}\,.$ (4.6)
In obtaining relations (4.6), the commutators
$[\mathcal{L}_{mn},e_{k}]$ $=$ $\eta_{nk}e_{m}-\eta_{mk}e_{n}\,,$ (4.7)
$[\mathcal{L}_{mn},\omega_{klr}]$ $=$ $\eta_{nk}\omega_{mlr}-\eta_{mk}\omega_{nlr}+\eta_{nl}\omega_{kmr}-\eta_{ml}\omega_{knr}+\eta_{nr}\omega_{klm}-\eta_{mr}\omega_{kln}\,.$ (4.8)
are used, as well as the equalities
$e_{(m}\omega_{n)}{}^{kl}=0\,,\qquad\omega_{(m}{}^{r[k}\omega_{n)r}{}^{l]}=0\,,$ (4.9)
that are valid for $\rm{AdS_{4}}$ space defined by the quantities (3.7)-(3.11).
Now, relations (4.6) together with the commutators
$[\mathcal{M}_{mn},(c\sigma_{l}\bar{c})]=\eta_{nl}(c\sigma_{m}\bar{c})-\eta_{ml}(c\sigma_{n}\bar{c})\,,\qquad[\mathcal{M}_{mn},(a\sigma_{l}\bar{a})]=\eta_{nl}(a\sigma_{m}\bar{a})-\eta_{ml}(a\sigma_{n}\bar{a})$ (4.10)
and $[P_{m},K]=[J_{mn},K]=0$ lead to the statement that the operators $L$ and $L^{+}$ defined in (2.6) and (2.7) are $\mathrm{SO}(2,3)$-invariant:
$[P_{n},L]=[P_{n},L^{+}]=0\,,\qquad[J_{mn},L]=[J_{mn},L^{+}]=0\,.$ (4.11)
In addition, using the commutation relations (4.6) and equalities (4.9) for the $\rm{AdS_{4}}$ metric and spin connection, we obtain that the operator $(D)^{2}$ defined in (2.12) commutes with all generators of the $\rm{AdS_{4}}$ algebra $\mathfrak{so}(2,3)$:
$[P_{m},(D)^{2}]=0\,,\qquad[J_{mn},(D)^{2}]=0\,.$ (4.12)
Note that the simplest way to prove the second relation in (4.12) is to use the following expression for the operator $(D)^{2}$:
$(D)^{2}=D_{m}D^{m}+[P_{m},D^{m}]\,.$ (4.13)
Thus, the conditions (4.12) lead to the statement that the operator $L_{0}$ defined in (2.5) is $\mathrm{SO}(2,3)$-invariant:
$[P_{m},L_{0}]=0\,,\quad[J_{mn},L_{0}]=0\,.$ (4.14)
In addition, it is easy to show that the operator (2.8) also commutes with the $\mathfrak{so}(2,3)$ algebra generators:
$[P_{m},U]=0\,,\quad[J_{mn},U]=0\,.$ (4.15)
Thus, the zero commutators (4.11), (4.14), (4.15) of all operators in the set (2.5)-(4.15) and all $\mathfrak{so}(2,3)$ generators (3.16), (3.17) lead to a very strong conclusion: the conditions (2.4) defining the infinite spin field in $AdS_{4}$ space are $\mathrm{SO}(2,3)$-invariant. As a result, we immediately face the question which representations of the $\mathrm{SO}(2,3)$ group correspond to the Fock space vectors satisfying the conditions (2.4). The next section is devoted to the discussion of this issue.
At the end of this section we introduce one more invariant operator of the algebra $\mathfrak{so}(2,3)$, which will be useful for subsequent consideration.
First, using the equality
$\mathcal{M}^{mn}\mathcal{M}_{mn}=M^{mn}M_{mn}+\bar{M}^{mn}\bar{M}_{mn}=-N(N+2)-\bar{N}(\bar{N}+2)\,,$ (4.16)
we can be represent the operator $L_{0}$ (2.5) in the form:
$L_{0}=(D)^{2}-\frac{1}{2R^{2}}\,\mathcal{M}_{mn}\mathcal{M}^{mn}+\frac{2\left(1+\boldsymbol{\mu}R\right)}{R^{2}}\,.$ (4.17)
Second, let us define the vector operator
$\Pi_{m}:=\mathcal{M}_{mn}D^{n}\,,$ (4.18)
which has the following commutation relations with the $\mathfrak{so}(2,3)$ generators:
$[P_{m},\Pi_{k}]=\omega_{mkl}\Pi^{l}\,,\qquad[J_{mn},\Pi_{k}]=\eta_{nk}\Pi_{m}-\eta_{mk}\Pi_{n}\,.$ (4.19)
Using these commutators and equalities (4.9) for the $\rm{AdS_{4}}$ spin connection, we obtain that the operator
$(\Pi)^{2}:=\Pi_{m}\Pi^{m}-\mathcal{M}_{mn}\omega^{mnl}\Pi_{l}$ (4.20)
commutes with all $\mathfrak{so}(2,3)$ generators:
$[P_{m},(\Pi)^{2}]=0\,,\qquad[J_{mn},(\Pi)^{2}]=0\,$ (4.21)
and hence it is an invariant of the $\mathfrak{so}(2,3)$ algebra.
5 Casimir operators and irreducible unitary representations
The $\mathfrak{so}(2,3)$ algebra is characterized by two Casimir operators. In this section we show that in Fock space of the vectors satisfying the conditions (2.4) the Casimir operators are expressed through the operator constraints (2.5)-(2.8) and describe the corresponding $\mathrm{SO}(2,3)$ irreducible representation.
5.1 Second order Casimir operator
The $\mathfrak{so}(2,3)$ algebra (3.3) (or the same (3.5)) has the following second order Casimir operator
$\mathcal{C}_{2}={}-\frac{1}{2}\,J^{MN}J_{MN}=R^{2}\,P^{m}P_{m}-\frac{1}{2}\,J^{mn}J_{mn}\,.$ (5.22)
Direct calculations lead to the following explicit expression for this operator
$P_{m}P^{m}-\frac{1}{2R^{2}}\,J^{mn}J_{mn}=(D)^{2}-\frac{1}{2R^{2}}\,\mathcal{M}^{mn}\mathcal{M}_{mn}\,.$ (5.23)
Using this relation together with expression (4.17), one gets for the Casimir operator (5.22)
$\mathcal{C}_{2}=R^{2}{L}_{0}-2\left(1+\boldsymbol{\mu}R\right)\,.$ (5.24)
Thus, the Casimir operator $\mathcal{C}_{2}$ is defined by the operator constraint $L_{0}$. Therefore, on the vectors (2.1) corresponding to the infinite spin field it acts as the operator of multiplication by a constant $-2(1+\boldsymbol{\mu}R)$.
5.2 Fourth order Casimir operator
The analysis of the fourth-order Casimir operator is not so simple as in the case of the second order Casimir operator.
The fourth order Casimir operator is defined as the square of the 5-vector
$W_{M}:=\frac{1}{8}\,\varepsilon_{MNKLI}J^{NK}J^{LI}\,.$ (5.25)
After some transformations, one gets (see e.g., [9])
$\mathcal{C}_{4}=W_{M}W^{M}=\frac{1}{2}\,\mathcal{C}_{2}\left(\mathcal{C}_{2}+3\right)-\frac{1}{4}\,J_{MN}J^{NK}J_{KL}J^{LM}\,,$ (5.26)
where $\mathcal{C}_{2}$ is defined in (5.22).
Let us now rewrite the last term in (5.26) using the generators $J_{mn}$ and $P_{m}$ defined in (3.4). We get
$J_{MN}J^{NK}J_{KL}J^{LM}$ $=$ ${}-4R^{2}\,P_{m}J^{mn}P^{k}J_{kn}+2R^{4}\,(P_{m}P^{m})^{2}+6R^{2}\,P_{n}P^{n}$ (5.27)
${}-4J^{mn}J_{mn}+J_{mn}J^{nk}J_{kl}J^{lm}\,.$
Relation (5.27) is transformed as follows. First, one expresses $P_{m}P^{m}$ through $\mathcal{C}_{2}$ using (5.22), then substitutes the result into (5.26) and uses all above relations in (5.27). Then one gets that the first term in (5.27) $\frac{1}{2}\,\mathcal{C}_{2}\left(\mathcal{C}_{2}+3\right)$ is completely canceled out. As a result, we obtain
$\mathcal{C}_{4}={}-\frac{1}{2}\,J^{mn}J_{mn}\,\mathcal{C}_{2}+R^{2}P_{m}J^{mn}P^{k}J_{kn}+\frac{1}{4}\,J^{mn}J_{mn}-\frac{1}{8}\,(J^{mn}J_{mn})^{2}-\frac{1}{4}\,J_{mn}J^{nk}J_{kl}J^{lm}\,.$ (5.28)
This operator can be represented in the form (see also a similar expression in [9]):
$\mathcal{C}_{4}={}-\frac{1}{2}\,\mathcal{M}^{mn}\mathcal{M}_{mn}\,\mathcal{C}_{2}+R^{2}(\Pi)^{2}-A(\mathcal{M})\,,$ (5.29)
where
$A(\mathcal{M}):=\frac{5}{4}\,\mathcal{M}^{mn}\mathcal{M}_{mn}+\frac{1}{8}\,(\mathcal{M}^{mn}\mathcal{M}_{mn})^{2}+\frac{1}{4}\,\mathcal{M}_{mn}\mathcal{M}^{nk}\mathcal{M}_{kl}\mathcal{M}^{lm}\,,$ (5.30)
and the $\mathfrak{so}(2,3)$-invariant operator $(\Pi)^{2}$ is given in (4.20).
Let us distinguish in the fourth-order Casimir operator, the operators (2.5)-(2.8). To do this, we represent the operator $(\Pi)^{2}$ (4.20) in the following form:
$(\Pi)^{2}=\mathcal{M}^{lm}\mathcal{M}_{l}{}^{n}\left(D_{m}D_{n}+\omega_{mnk}D^{k}\right)\,.$ (5.31)
After using equalities (4.16),
$\mathcal{M}_{mn}\mathcal{M}^{nl}\mathcal{M}_{lr}\mathcal{M}^{rm}=-\mathcal{M}^{mn}\mathcal{M}_{mn}+\frac{1}{4}\,(\mathcal{M}^{mn}\mathcal{M}_{mn})^{2}+N(N+2)\bar{N}(\bar{N}+2)\,,$ (5.32)
$\mathcal{M}^{l(m}\mathcal{M}_{l}{}^{n)}={}S(N,\bar{N})\,\eta^{mn}-(c\sigma^{(m}\bar{c})(a\sigma^{n)}\bar{a})\,,$ (5.33)
where
$S(N,\bar{N}):={}-\frac{1}{4}\,\Big{[}N(N+2)+\bar{N}(\bar{N}+2)-2N\bar{N}\Big{]}\,,$ (5.34)
we get that in the $\rm{AdS_{4}}$ case with relations (3.8)-(3.11) the operator (5.31) takes the form:
$(\Pi)^{2}$ $=$ $S(N,\bar{N})L_{0}+L^{+}L+\left(\boldsymbol{\mu}+\frac{K^{2}}{4R}\right)L+\left(\boldsymbol{\mu}+\frac{(K-2)^{2}}{4R}\right)L^{+}$
${}+\frac{1}{R^{2}}\,A(\mathcal{M})+\frac{1}{R^{2}}\,(1+\boldsymbol{\mu}R)\bigl{(}\boldsymbol{\mu}R-\mathcal{M}_{mn}\mathcal{M}^{mn}\bigl{)}\,,$
where $S(N,\bar{N})$ and $A(\mathcal{M})$ are defined in (5.30) and (5.34), respectively. As a result, the fourth order Casimir operator (5.29) is written as
$\mathcal{C}_{4}={}R^{2}\left[\frac{1}{4}K(K-2)L_{0}+L^{+}L+\left(\boldsymbol{\mu}+\frac{K^{2}}{4R}\right)L+\left(\boldsymbol{\mu}+\frac{(K-2)^{2}}{4R}\right)L^{+}\right]+\boldsymbol{\mu}R(1+\boldsymbol{\mu}R)\,.$ (5.36)
One can prove that the operator (5.36) is Hermitian. We see that the operator $\mathcal{C}_{4}$, as well as the operator $\mathcal{C}_{2}$, is expressed through the operator constraints (2.5)-(2.7).
5.3 Irreducible representation
Using the expressions (5.24) and (5.36), we can determine the action of the $\mathfrak{so}(2,3)$ Casimir operators on the infinite spin vectors (2.1) in $\rm{AdS_{4}}$ space satisfying the conditions (2.4).
From the first condition $L_{0}|{\Psi}\rangle=0$ of the set (2.4) we obtain that the infinite spin vector (2.1) is the eigenvector of the second-order Casimir operator
$\mathcal{C}_{2}|{\Psi}\rangle=c_{2}|{\Psi}\rangle\,$ (5.37)
with the eigenvalue
$c_{2}=-2(1+\boldsymbol{\mu}R)\,.$ (5.38)
Besides, due to the first three conditions in (2.4), namely, the conditions $L_{0}|{\Psi}\rangle=0$, $L|{\Psi}\rangle=0$, $L^{+}|{\Psi}\rangle=0$, the same vector $|{\Psi}\rangle$ is also the eigenvector of the fourth order Casimir operator
$\mathcal{C}_{4}|{\Psi}\rangle=c_{4}|{\Psi}\rangle$ (5.39)
with the eigenvalue
$c_{4}=\boldsymbol{\mu}R(1+\boldsymbol{\mu}R)\,.$ (5.40)
Thus, the vectors (2.1) under the conditions (2.4) are the eigenvectors of all Casimir operators of the $\mathfrak{so}(2,3)$ algebra with the continuous eigenvalues (5.38) and (5.40). Therefore, the vector $|{\Psi}\rangle$ describing the infinite spin field dynamics in $\rm{AdS_{4}}$ space corresponds to the continuous irreducible representation of the $\mathrm{SO}(2,3)$ group. The last condition in (2.4) removes the degeneracy of the spectrum associated with the use of additional spinor operators.
Relations (5.38) and (5.40) show that the eigenvalues of the Casimir operators are related to each other by the relation
$c_{4}=\frac{c_{2}}{2}\left(\frac{c_{2}}{2}+1\right)$ (5.41)
and are characterized by one dimensionless continuous real parameter $\boldsymbol{\mu}R$. This means that of the two Casimir operators, only one is independent
$\mathcal{C}_{4}=\frac{1}{2}\,\mathcal{C}_{2}\left(\frac{1}{2}\,\mathcal{C}_{2}+1\right).$ (5.42)
Such representations, where of all Casimir operators only one is independent, are called the most degenerate representations in the mathematical literature. Therefore, the Fock space vector describing the Lagrangian dynamics of the infinite spin field in $\rm{AdS_{4}}$ space corresponds to the most degenerate continuous unitary representation of the $\rm{AdS_{4}}$ symmetry group.
In the paper [9], using the light-cone formalism, Metsaev proposed a classification of classically unitary infinite spin fields in $\rm{AdS}$ space. The classification is based on relations (3.19) and (3.20)-(3.24) from [9] for complex coefficients $p$ and $q$ in terms of which the eigenvalues of the second- and fourth-order Casimir operators in $\rm{AdS}_{4}$ space take the form [9]:
$c_{2}=p^{2}+q^{2}-\frac{5}{2}\,,\qquad c_{4}=\left(p^{2}-\frac{1}{4}\right)\left(q^{2}-\frac{1}{4}\right),$ (5.43)
which are symmetric with respect to replacements $p\leftrightarrow q$. We see that our obtained results (5.38), (5.40) correspond to
$p=\varepsilon_{p}\left(\frac{1}{2}+i\sqrt{\boldsymbol{\mu}R}\right),\qquad q=\varepsilon_{q}\left(\frac{1}{2}-i\sqrt{\boldsymbol{\mu}R}\right),$ (5.44)
where $\varepsilon_{p}=\pm 1$ and $\varepsilon_{q}=\pm 1$ independently of each other. In the case $\boldsymbol{\mu}>0$, the values (5.44) obtained are included in the set of solutions $\mathbf{ii}$ and $\mathbf{iii}$ from [9], whereas (5.44) are the solutions from the set $\mathbf{vi}$ in the case $\boldsymbol{\mu}<0$, that is, $p$ and $q$ become real. However, the cases $\mathbf{ii}$, $\mathbf{iii}$, $\mathbf{vi}$ as well as the other cases in [9] mean that two Casimir operators $\mathcal{C}_{2}$ and $\mathcal{C}_{4}$ are independent while in our case only one of them is independent. As we have already pointed out, our case corresponds to the most degenerate representation. Perhaps this degeneration is due to a special form of operator constraints (2.5), (2.6), (2.7).
6 Summary
In this paper we have studied the algebraic aspects of the Lagrangian formulation for the infinite spin field model proposed in [12]. The formulation is constructed in the framework of the BRST approach to higher spin field theory. The main elements of this approach are the Fock space vectors $|{\Psi}\rangle$ (2.1) and the operator constraints (2.5)-(2.8) in this space, forming a closed algebra in terms of commutators, which ensures gauge invariance of the Lagrangian theory. The operator constraints in the case under consideration includes the radius $R$ of $\rm{AdS}_{4}$ space and the arbitrary real massive parameter $\boldsymbol{\mu}$ associated with the continuous representation. It was proved that the above operator constraints are the invariants of the $\mathrm{SO}(2,3)$ group and the Casimir operators of the $\mathfrak{so}(2,3)$ algebra in the Fock space are completely expressed through these constraints in the form (5.24) and (5.36).
It turned out that the Casimir operators are not independent but are related to each other by (5.42). Since there is only one independent Casimir operator in the given representation, we conclude that the representation under consideration is most degenerate. One can show that the constructed representations correspond to the cases $\mathbf{ii}$, $\mathbf{iii}$ (when $\boldsymbol{\mu}>0$) and $\mathbf{vi}$ (when $\boldsymbol{\mu}<0$) in the Metsaev classification of the unitary continuous spin irreducible representations in $\rm{AdS}$ space [9].
Acknowledgments
Authors are grateful to R.R. Metsaev for stimulating discussions and comments. The authors are also thankful to M.A. Vasiliev and Yu.M. Zinoviev for useful comments. M.A.P. is grateful to A.A. Golubtsova for helpful discussions. The work of M.A.P. was supported by the JINR AYSS Foundation, Project No 24-301-03.